OpenStudy (anonymous):

How do you write r=2sintheta and r=2sin2theta into polar coordinates?

OpenStudy (anonymous):

it is polar coordinates

OpenStudy (anonymous):

r = 2 sin theta, theta is theta

OpenStudy (anonymous):

maybe you mean cartesian coordinates?

OpenStudy (anonymous):

it says solve the system of equations on the polar plane: then it has those two equations. And says Write your solutions as polar coordinates [hint there are 4]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

yeah ha. Do you know anyone who could help?

OpenStudy (anonymous):

yes me, give me medal

OpenStudy (anonymous):

I don't know how to?

OpenStudy (anonymous):

i just gave you medal

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

i took it away your medal

OpenStudy (anonymous):

solve 2 sin t = 2 sin (2t)

OpenStudy (anonymous):

Ah, it's already in polar coordinates. (r, theta)

OpenStudy (anonymous):

assume that r1 = r2 , where r1 = 2 sin t , and r2 = 2 sin (2t)

OpenStudy (anonymous):

2 sin t = 2 ( 2 sin t cos t )

OpenStudy (anonymous):

woah slow down haha

OpenStudy (anonymous):

2sin t - 4sint*cost = 0

OpenStudy (anonymous):

more medals

OpenStudy (anonymous):

2sin t ( 1 - 2 cos t ) = 0

OpenStudy (anonymous):

how is that four solutions

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

solve what?

OpenStudy (anonymous):

2sin t = 0 or 1- cos t = 0

OpenStudy (anonymous):

helppppp

OpenStudy (anonymous):

sin t = 0 when t = 0, pi ,

OpenStudy (anonymous):

1- 2 cos t = 0 when cos t = 1/2 , when t = pi/3, 5pi/3

OpenStudy (anonymous):

so the solutions are t = { 0, pi, pi/3, 5pi/3}

OpenStudy (anonymous):

how do i show the work? what do I write for the equations and what answer goes with which one

OpenStudy (anonymous):

Did you use the right equations? 2sint and 2sin2t

OpenStudy (anonymous):

never mind i got it. thanks (: